The ratio of the weight of copper to the weight of zinc in a coin is 1/38. If 55 kg of zinc is being turned into coins, how much copper is needed?
Answer by checkley79(3341)

1/38=x/55 cross multiply.
38x=55
x=55/38
x=1.48 oz. of copper is needed.